OpenStudy (anonymous):

2y-3+(4x+8)i=0 help me find x and y

OpenStudy (aaronandyson):

Is it (4x +0)y rather??

OpenStudy (anonymous):

nope

OpenStudy (michele_laino):

please you have to solve these equation separately: \[\begin{gathered} 2y - 3 = 0 \hfill \\ 4x + 8 = 0 \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

y=3/2 and x=-2??

OpenStudy (michele_laino):

yes!

OpenStudy (aaronandyson):

Even I got that It's correct.

OpenStudy (michele_laino):

more explanation: you have to set these conditions: real part =0 imaginary part =0

OpenStudy (anonymous):

thanks! i was just making sure of my ans :)

OpenStudy (michele_laino):

thanks! :)